1) Before afternoon calibrations, I put the PropID in the GUI (newfirm_default) sometime during the night this was erased. I don't know if this happened because of a software glitch or if the GUI were restarted while I was't looking. Not all of the data from the night may have made it to the archive, so I ask that Sean McManus look to see what files were staged tonight.
(2) I remember at least two images had problems with syncing after work on the guider was completed. One ocurred shortly after the telescope was handed over to me (00:30 local). The other occurred when I was taking out of focus Ks band images to check the position/alignment of the Lyot stop. Even though there were at least two sync issues, this was a great improvement over the Aprill engineering time. I will work with Marco and Peter to resolve this.
(3) Hernan wrote JIRA for ITOps. Some of the guider software displays that Simon created can not be opened on the operators' console. Bernardo is coming to Blanco tomorrow to resolve this. Javier has some (all?) of the details.
Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
---|---|---|---|---|---|---|
2023-07-04 | 0.0 | 11.25 | 0.0 | 0.0 | 0.0 | 11.25 |
Links
[1] mailto:sean.points@noirlab.edu
[2] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/42673/delete?destination=print/content/enr-ctio-blanco-4-m-2023-07-04
[3] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/42673/edit?destination=print/content/enr-ctio-blanco-4-m-2023-07-04
[4] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/add/node/20619?destination=print/content/enr-ctio-blanco-4-m-2023-07-04
Work done:
(1) Installation of new 20W lamps for dome flats. Because of this, I wasn't able to use the lamp intensity and exposure times I determined for the filters in April. The counts are still a little bit high in some filters (> 5000). I will spend time tomorrow trying to refine the default setups.
(2) Worked on the zero-point (telescope pointing procedure). We imaged a bright star in the upper-left corner of detector 3 and then detemined an offset to the center of the detector gap. Prior documentation states that the detector gap is 35". I think this is incorrect. Tests from astrometrically reduced data indicated that the detector gap is 50". Results from tonight may indicate that the gap is 70". More tests will be performed tomorrow. The pointing and default astrometry of the detectors is pretty good.
(3) After zero-point, we performed a telescope focus using the J filter. After fiding the best focus for NEWFIRM, Rolo, Omar, and Marco began the task of finding the best focus for the north guider. This was found and is fairly close to the middle of the guider focus range. Great job by Ryan, Andres, Rolo, Omar, and Marco for correctly determining the proper solution after the April run. We also checked offsets with the guider and found the North guider camera to have an orientation of East left and North down. The rotation angle of the field was also determined to be 0.4deg.
(4) Guider tests were performed using the South guider. The orientation is the same - East left and North down. The roatation angle of South camera was calculated to be 7.0deg.
(5) We performed telescope focus sequences for JHKs to determine the focus offsets among the filters. The values are similar to those determined in April. Rolo, Omar, and Marco verified that guider could still focus for these filters. The documenation for NEWFIRM from the Mayall at KPNO states that the largest focus difference between filters is the offset between J and Ks. If this holds true, we should have no trouble with telescope focus and guider focus during observations.
(6) After work with the guider was completed for the night, I took a series of JHKs observations of Persson 9181 with different exxposure times and coadds. These data need to be analyzed . First glance has me worried because the counts/flux recorded in a 2sec exposure is more than 2x the counts/flux recorded in a 1sec exposure. I also note that the ratios between exposures of 5sec (5sec 1 coadd), 30sec (5sec 6 coadd) and 60sec (5sec 12 coadd) seem to be normal. This is a priority.
All seeing values from RINGSS
All in all, a very good and productive night. Many thanks to everybody who have worked hard on the NEWFIRM project over the past three years to get us to this stage. As others have mentioned today, we are very close to being ready for visiting astronomers to use NEWFIRM.